notmeta:

(algebra 1) need help with a few questions

notmeta:

(algebra 1) n

Chloe83:

The equation you're looking for is y-y1=m(x-x1). So you're slow is outside the parentheses, your x is inside, and your y is one the left side of the equal sign

AcornTheNut:

@toga

toga:

@acornthenut wrote:
@toga
the question has been answered so there's no need to tag me : )

AcornTheNut:

@toga wrote:
@acornthenut wrote:
@toga
the question has been answered so there's no need to tag me : )
alright

notmeta:

@toga wrote:
@acornthenut wrote:
@toga
the question has been answered so there's no need to tag me : )
it hasn't. i don't understand what he said

toga:

@notmeta wrote:
@toga wrote:
@acornthenut wrote:
@toga
the question has been answered so there's no need to tag me : )
it hasn't. i don't understand what he said
ok then I will answer it

toga:

To find the equation in point-slope form that passes through the point (4, -2) with a slope of \( \frac{1}{2} \), we use the point-slope formula: \[ y - y_1 = m(x - x_1) \] Where \( (x_1, y_1) \) is the point (4, -2) and \( m \) is the slope \( \frac{1}{2} \). Plugging in the values: \[ y - (-2) = \frac{1}{2}(x - 4) \] This simplifies to: \[ y + 2 = \frac{1}{2}(x - 4) \] Thus, the equation in point-slope form is: C. \( y + 2 = \frac{1}{2}(x - 4) \) So, the correct answer is C.

toga:

Do you understand this better?

notmeta:

@toga wrote:
Do you understand this better?
way better thank you. I wasn't sure whether or not to add the y+y1 or to substract so this really helped. also i have a different problem that I need help with should I keep this post open or make a new one? also thank you again for the help

toga:

@notmeta wrote:
@toga wrote:
Do you understand this better?
way better thank you. I wasn't sure whether or not to add the y+y1 or to substract so this really helped. also i have a different problem that I need help with should I keep this post open or make a new one? also thank you again for the help
make a new one you're not actually supposed to post multiple questions in one post

toga:

But no one really cares